OpenStudy (anonymous):

Jesse rents a moving van for $75 and must pay $2 per mile. The following week, Alex rents the same van, is charged $80 for the rental and $1.50 per mile. If they each paid the same amount and drove the same number of miles, how far did they each travel?

OpenStudy (anonymous):

number of miles=x 75+2x=80+1.5x solve it

OpenStudy (anonymous):

so the answer would be .1

OpenStudy (anonymous):

no, 10. try again

OpenStudy (anonymous):

i minused 75 fist to each side and then got 2x =5+1.5 then minused 1.5 from each and got .5x=5 and divided x=.1

OpenStudy (anonymous):

the last bit was wrong. 0.5x=5 time 2 to both sides

OpenStudy (anonymous):

0.5*2=1

OpenStudy (anonymous):

so .5 is actually 5?

OpenStudy (anonymous):

0.5 is actually 1/2 so 1/2 times 2 = 1

OpenStudy (anonymous):

where does .5 come from

OpenStudy (anonymous):

2x-1.5x of course

OpenStudy (anonymous):

is that the first step

OpenStudy (anonymous):

that's nearly the last step. look at your working